Aerospace industry believes that the cost of a space project is a function of the weight of the major object being space sent into space. The following data represent weight in tons and the cost in $ millions.


Weights cost

1.59 54

3.20 150

0.56 20

0.89 25

1.45 40

1.90 120

2.57 124


calculate:

1) construct regression equation

2) What is the strength of the regression equation?


1
Expert's answer
2021-02-19T13:58:19-0500

1)

The regression equation:

"y=ax+b"

"a=\frac{(\sum y)(\sum x^2)-(\sum x)(\sum xy)}{n(\sum x^2)-(\sum x)^2}"

"a=\frac{533\cdot26.18-12.16\cdot1203.99}{7\cdot26.18-12.16^2}=-19.4"


"b=\frac{n(\sum xy)-(\sum x)(\sum y)}{n(\sum x^2)-(\sum x)^2}"

"b=\frac{7\cdot1203.99-12.16\cdot533}{7\cdot26.18-12.16^2}=55"


"y=-19.4x+55"



2)

Correlation coefficient:

"r=\frac{cov(x,y)}{\sqrt{s_x^2s_y^2}}"

"cov(x,y)=\frac{\sum (X-\overline{X})(Y-\overline{Y})}{n-1}"

"s_x^2=\frac{\sum (X-\overline{X})^2}{n-1}"

"s_y^2=\frac{\sum (Y-\overline{Y})^2}{n-1}"


"\overline{X}=\frac{\sum x}{n}=\frac{12.16}{7}=1.74"


"s_x^2=\frac{5.06}{6}=0.84"


"\overline{Y}=\frac{\sum y}{n}=\frac{533}{7}=76.14"


"s_y^2=\frac{17232.86}{6}=2872.14"


"cov(x,y)=278.1\/6=46.35"


"r=\frac{46.35}{\sqrt{0.84\cdot2872.14}}=0.94"


The regression has a strong positive correlation.